solve
4x+4y=3
-10+5y=-5x

Answer by Fermat(136)

4x+4y=3
-10+5y=-5x
rearrange as follows,
4(x+y) = 3
5x + 5y = 10 (can you see how the 2nd eqn can be rearranged to give this line?)
then,
x+y = 3/4 (divde the 1st eqn by 4 on both sides)
x+y = 2 (divde the 2nd eqn by 5 on both sides)
Hmmm,
You can't have x+y equal 3/4 and equal 2 at the same time!!
This means that the equations are inconsistent, and have no solution.
You didn't make a typo somewhere ?